OpenStudy (anonymous):

Find the x and y intercepts of 5x^2 + 20x

OpenStudy (ahsome):

The equation is: \(y=5x^2+20x\) Remember, in the \(y-intercept\), the \(x\) value is 0 In the \(x-intercept\), the \(y\) value is 0 To find the \(y-intercept\), substitute 0 for \(x\) \(y=5x^2+20x\) \(y=5*0^2+20*0\) \(y=5*0+0\) \(y=0\) To find the \(x-intercept\), substitute 0 for \(y\) \(y=5x^2+20x\) \(0=5*x^2+20x\) Then, you can either factorise, or use the quadratic equation: \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] Sub in the values: \[x=\frac{-20\pm\sqrt{20^2-4*5*0}}{2*5}\] That will give you the two \(x-intercepts\)

OpenStudy (ahsome):

Do you understand, @Dholmes44?